OpenStudy (anonymous):

quick math help ? Find the derivative of f(x) = -12x2 + 9x at x = 6. -112.5 -135 -90 -108

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x) = -12x^2 + 9x }\) apply the power rule to each term

OpenStudy (anonymous):

f(x)=24x+9x ??

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f{\tiny~~}'(x) = -(\color{red}{2})12x^{\color{red}{2}-1} + (\color{red}{1})9x^{\color{red}{1}-1} }\) \(\large\color{black}{ \displaystyle f{\tiny~~}'(x) = -24x + 9 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f{\tiny~~}'(\color{blue}{6}) =-24(\color{blue}{6})+9 }\)

OpenStudy (solomonzelman):

(all you got to do, is to simplify this)

OpenStudy (anonymous):

-135?